# Mixed Strategy Nash Equilibrium for Crowd Navigation

### v4

@@ V-B Multi-agent navigation evaluation
We select two baselines for comparison: (1) Optimal reciprocal collision avoidance (ORCA[8]; and (2) collision avoidance with deep reinforcement learning (CADRL[51]. We choose the selected baselines partially based on their widely accessible implementations. Furthermore, we choose ORCA as it is the de facto non-learning decentralized collision avoidance model for simulating crowds. We choose CADRL as it is the de facto learning-based crowd navigation model.
@@
TABLE III: Computation time of game-theoretic methods.
@@ Results
The evaluation results are presented in Table I(b) (safety) and Table II (navigation efficiency). Note that ORCA guarantees collision-free paths, thus it has $0\%$ collision rate across all experiments. On the other hand, our method BRNE generates joint navigation strategies with larger safety distance, better efficiency, and lower disparity. Even though BRNE strategies are not collision-free, the collision rate is still minimal. Since the multi-agent navigation benchmark assumes all agents follow the same behavioral model, having a $0\%$ collision rate in this benchmark does not necessarily mean collision-free with humans. More importantly, as shown in Figure 5, the collision-free guarantee of ORCA could significantly compromise the path quality, leading to highly unnatural joint strategies when agents are unevenly distributed. This indicates that, when used as a coupled prediction and planning framework on the robot, ORCA could plan the robot’s path based on highly unrealistic expectations of human reactions. Lastly, the learning-based baseline CADRL performs poorly across all metrics and exhibits highly non-smooth trajectories as shown in Figure 5. Note that the CADRL policy is trained using the toolbox provided in [52] with pedestrians simulated with ORCA. Thus, the trained CADRL policy is a collision avoidance policy in crowds. We test the CADRL policy in this multi-agent navigation benchmark to verify if the policy is a coupled prediction and planning method, instead of a dynamic object avoidance method. Videos of the multi-agent navigation experiments are included in our project website: https://sites.google.com/view/brne-crowdnav.
TABLE IV: Results of simulated crowd navigation experiments.
@@
We use the multi-agent navigation benchmark to evaluate how the number of agents and samples affect the computation time of BRNE. We fixed the planning horizon of BRNE at 50 time steps (5 seconds), then we vary the number of agents from 4 to 8 and the number of samples per agent from 100 to 500. The tested BRNE implementation uses Numba for parallelization and runs on a laptop with Intel Core i9-12900K Processor (20 threads). The results are shown in Figure 7. We can see that when the algorithm is accelerated through parallelization in practice, the computation complexity is close to linear with respect to the number of agents. Since the scalability with the number of agents is often a computation bottleneck for game-theoretic planners, in Table III, we further compare the computation speed of BRNE with two state-of-the-art dynamic game solvers, ALGAMES [60] and iLQGames [59]. We use 100 samples per agent for BRNE and use the results reported in the “intersection” experiment from [60], as it is the most similar to the crowd navigation setting in this test. We use the importance-sampling-based BRNE implementation throughout this test. In Table III, we can see that BRNE is at least one order of magnitude faster than ALGAMES and iLQGames and is the only method with sufficient computation time for real-time planning with 4 agents. ALGAMES and iLQGames are general-purpose game-theoretic planners that can be applied to applications other than crowd navigation, while BRNE is designed specifically for crowd navigation. BRNE explicitly benefits from its narrow scope for better computation efficiency in dense crowds.
@@
V-C Simulated crowd navigation evaluation
@@
TABLE V: Results of human dataset crowd navigation experiments.
@@
We select $3$ baselines to compare with: (1) ORCA; (2) CADRL; (3) Constrastive learning social navigation (NCE[54]. We choose ORCA and CADRL for the same reason as we choose them for multi-agent navigation evaluation. We choose NCE since it is reported as the state-of-the-art method in simulated crowd navigation benchmarks. We train NCE in the environment provided by [54].
@@ Results
In Table IV and Figure 8, we show the safety and navigation efficiency results. In Figure 9, we show representative frames of our method from the simulated tests. We can see that our method BRNE is competitive on both navigation safety and efficiency with the state-of-the-art reinforcement learning method NCE. Both NCE and CADRL are tested in the same simulation environment where it is trained, while our method BRNE requires no training. CADRL performs poorly in this benchmark, exhibiting a high collision rate and low navigation efficiency. Furthermore, CADRL fails to reach the navigation goal and we have to terminate the trial after a certain amount of time, causing the low standard deviation in the time-to-goal metric. The performance of CADRL could be due to the difference between the testing environment provided by [54] and the training environment of CADRL. Videos of the simulated crowd navigation experiments are included in our project website: https://sites.google.com/view/brne-crowdnav.
V-D Crowd navigation in human datasets
@@
We evaluate crowd navigation with prerecorded human pedestrian behaviors within the same geometrical spaces of recordings. The evaluation is conducted in SocNavBench [81], a state-of-the-art, high-fidelity crowd navigation benchmark framework. SocNavBench contains human pedestrian data from two of the most commonly used pedestrian datasets ETH [82] and UCY [83]. A set of 33 curated episodes are extracted from the original datasets for benchmarking, which assemble a set of highly interactive test trials. The curated episodes contain an average of 44 pedestrians per episode. The distance between the goal location and the start location for the robot ensures that the navigation task can be finished within 25 seconds if the robot does not avoid collision and navigates with the maximum permitted velocity of $1.2m/s$. An occupancy grid map of the real-world space is provided for each episode, the robot has access to the map a priori and can only navigate in the free space. Note that, even though the pedestrian agents in SocNavBench are non-reactive given the pre-recorded human behavior, the robot agent still assumes the pedestrian agents are real-world reactive humans. We refer readers to [81] for more details about the benchmark, as well as the discussion on the choice of using pre-recorded human behavior. In this experiment, BRNE is implemented with importance sampling as a model predictive planner, updating its inference of mixed strategy Nash equilibrium based on the latest observation of other agents’ positions and velocities.
@@
We compare our method BRNE with the four baselines and the corresponding results reported in the original paper of SocNavBench [81]. They are (1) the social force model (SF[7]; (2) ORCA; (3) CADRL; and (4) a pedestrian-unaware baseline provided by the benchmark with static obstacle avoidance capability (Baseline). Note that ORCA and CADRL do not have static obstacle avoidance capability. Thus, they use Baseline as a meta-planner to avoid static obstacles in the benchmark. We tried to use Baseline directly as the meta-planner to generate the nominal mixed strategy for BRNE. However, even though Baseline is sufficient for BRNE to avoid static obstacles, it is not sufficient for BRNE to consistently reach the goal—recall that BRNE requires a strong goal-oriented nominal mixed strategy. Therefore, we slightly modified Baseline to keep its static obstacle avoidance module but improve its goal-reaching capability. We name this modified baseline planner Meta-Planner and report its performance as well. We will release the implementation of both BRNE and Meta-Planner in this benchmark.
@@ Results
Table V shows results of the human dataset crowd navigation experiments. Figure 10 shows representative snapshots of the benchmark tests. Both our method (BRNE) and the social force model (SF) significantly outperform the rest of the methods on navigation safety. Our method shares the minimal number of collisions with the social force model while being the only method with zero freezing behavior and exhibiting better path efficiency compared with the social force model. More importantly, compared to our meta-planner baseline Meta-planner(Ours), whose path is used by BRNE to generate the nominal mixed strategy, BRNE significantly improves navigation safety, reducing the number of pedestrian collisions from 37 to 1, while sharing nearly identical navigation efficiency as the baseline. As discussed in the next subsection, this one collision incident of our method is due to the pre-processing of the dataset instead of our method’s performance. These experimental results show that, when interacting with real human crowds in unstructured environments, our method can improve an existing navigation method’s safety to be near-perfect with minimal impact on navigation efficiency. This indicates that our method reaches human-level crowd navigation performance with a meta-planner on safety and navigation efficiency. The results also show great potential for deploying our method on other navigation pipelines. Videos of the human dataset experiments are included in our project website: https://sites.google.com/view/brne-crowdnav.
@@
While we will demonstrate our BRNE algorithm on robot hardware with real-world humans in the next section, we leave benchmarking our method in real-world crowds for future work, due to the time and financial cost of conducting such benchmarks. Here, we want to discuss the potential gap of sim-to-real transfer based on the human dataset benchmark results. First, since the robot is observing real-world human behavior in the benchmark, the benchmark results can at least verify which method will fail in the real world—a method that is dangerous or inefficient in the benchmark will likely remain dangerous or inefficient in the real world. Second, one main challenge of sim-to-real transfer in crowd navigation is the misalignment of human behavior assumptions. In our case, by inferring mixed strategy Nash equilibrium for navigation, the robot assumes the humans are cooperative, which is not necessarily true. However, the human dataset benchmark can serve as a preliminary verification of the influence of the misaligned human behavior assumption. Recall that, even though human agents are not reactive in the benchmark, the robot still believes it is interacting with real humans and observing real human behavior. The numerical efficiency of our method enables fast replanning during navigation, allowing the robot to constantly adjust its plan based on the latest observation, including observations that humans do not follow the assumed cooperative behavior. Furthermore, the robot makes navigation decisions based on mixed strategies—it does not anticipate exact human behavior but instead a distribution of likely human behaviors, which already considers unexpected human reactions. Thus, the fast replanning capability and mixed strategy-based modeling make the performance of our method in the human dataset benchmark more transferable in the real world. Lastly, our method is essentially built on top of an existing meta-planner provided by the benchmark. The modular design of our algorithm enhances the deployment flexibility, which lowers the gap of sim-to-real transfer. For example, our method can be integrated into existing robot navigation frameworks such as the navigation stack from the Robot Operating System [84, 85]. This is particularly important for crowd navigation in unstructured real-world environments, which most existing crowd navigation methods and all game-theoretic planning methods do not address. Lastly, even though out of the scope of this paper, we want to point out one important aspect of the sim-to-transfer that is not addressed in the benchmark: the perception aspect of crowd navigation. The benchmark assumes near-perfect robot perception of pedestrian states, but robust perception for both robot localization and pedestrian tracking remains an open challenge and could influence the real-world deployment of our method.
V-E Real-World Hardware Demonstration
@@
We implement our BRNE algorithm within the Isaac Robot Operation System (ROS) framework to communicate with the perception module and the robot’s low-level controller. We implement the algorithm in Python and used the PyTorch library to accelerate the weight updating step of the algorithm using the low-power GPU of the onboard computer. To maintain the real-time computation speed, we limited the maximum number of agents processed by the algorithm to 5, which allows the algorithm to control the robot at 10 Hz and replan at 5 Hz, with a planning horizon of 2 seconds and 200 trajectory samples per agent. We will release the hardware design and software implementation alongside the algorithm parameters.
@@
The demonstrations are conducted in a $3m$ wide and $9m$ long space inside an indoor atrium. The demonstrations are designed to mimic the specifications of the crowd navigation tests with ORCA agents (such as the ones shown in Figure 9) and with pedestrian datasets (such as the ones shown in Figure 10). The number of pedestrians varies from $4$ to $10$. Our demonstrations follow common design principles in other crowd navigation works, such as [43] and [45]. To mimic the test with ORCA agents, pedestrians were instructed to go to the designated goals in a circle. To mimic the pedestrian dataset benchmark, pedestrians were instructed to move freely in the space. Across all demonstrations, the pedestrians are instructed to “move with normal walking speed and treat the robot as a walking person or dog”. The navigation task for the robot is to move safely and efficiently from one side of the room to the other, with a traveling distance varying between $6m$ to $8m$. The robot moves at a nominal speed of $0.5m/s$.
VI Conclusion
@@
In this work, we propose a computation-efficient mixed strategy Nash equilibrium model for crowd navigation. Mixed strategy Nash equilibrium provides a high-level cooperation model for the robot to plan actions that leverage human cooperation for collision avoidance while maintaining the uncertainty in human behavior. Despite the general hardness of computing mixed strategy Nash equilibrium, we achieve real-time inference speed on a laptop CPU and a low-power embedded computer by establishing the formal connection between mixed strategy Nash equilibrium and a simple iterative Bayesian update scheme. We name the proposed model Bayes’ Rule Nash Equilibrium (BRNE). We further develop a model predictive crowd navigation framework using Gaussian processes to bootstrap agents’ nominal mixed strategies. The proposed model can be incorporated into existing navigation frameworks to navigate alongside humans in unstructured environments with static obstacles. Our experiment results show that our BRNE model significantly improves the safety of a human-unaware planner without compromising navigation efficiency in the human dataset benchmark, reaching human-level performance. Compared to other crowd navigation algorithms, our model consistently outperforms them in both safety and navigation efficiency. Lastly, we demonstrate the practicality of our model on an untethered robot quadruped for real-time crowd navigation with fully onboard perception and computation.

### v6

@@ 5.2 Multi-agent navigation evaluation
We select two baselines for comparison: (1) Optimal reciprocal collision avoidance (ORCA(van den Berg et al., 2011); and (2) collision avoidance with deep reinforcement learning (CADRL(Chen et al., 2017). We choose the selected baselines partially based on their widely accessible implementations. Furthermore, we choose ORCA as it is the de facto non-learning decentralized collision avoidance model for simulating crowds. We choose CADRL as it is the de facto learning-based crowd navigation model.
@@
Table 3: Computation time of game-theoretic methods.
@@ Results
The evaluation results are presented in Table 1(b) (safety) and Table 2 (navigation efficiency). Note that ORCA guarantees collision-free paths, thus it has $0\%$ collision rate across all experiments. On the other hand, our method BRNE generates joint navigation strategies with larger safety distance, better efficiency, and lower disparity. Even though BRNE strategies are not collision-free, the collision rate is still minimal. Since the multi-agent navigation benchmark assumes all agents follow the same behavioral model, having a $0\%$ collision rate in this benchmark does not necessarily mean collision-free with humans. More importantly, as shown in Figure 5, the collision-free guarantee of ORCA could significantly compromise the path quality, leading to highly unnatural joint strategies when agents are unevenly distributed. This indicates that, when used as a coupled prediction and planning framework on the robot, ORCA could plan the robot’s path based on highly unrealistic expectations of human reactions. Lastly, the learning-based baseline CADRL performs poorly across all metrics and exhibits highly non-smooth trajectories as shown in Figure 5. Note that the CADRL policy is trained using the toolbox provided in Chen et al. (2019) with pedestrians simulated with ORCA. Thus, the trained CADRL policy is a collision avoidance policy in crowds. We test the CADRL policy in this multi-agent navigation benchmark to verify if the policy is a coupled prediction and planning method, instead of a dynamic object avoidance method. Videos of the multi-agent navigation experiments are included in our project website: https://sites.google.com/view/brne-crowdnav.
Table 4: Results of simulated crowd navigation experiments.
@@
We use the multi-agent navigation benchmark to evaluate how the number of agents and samples affect the computation time of BRNE. We fixed the planning horizon of BRNE at 50 time steps (5 seconds), then we vary the number of agents from 4 to 8 and the number of samples per agent from 100 to 500. The tested BRNE implementation uses Numba for parallelization and runs on a laptop with Intel Core i9-12900K Processor (20 threads). The results are shown in Figure 7. We can see that when the algorithm is accelerated through parallelization in practice, the computation complexity is close to linear with respect to the number of agents. Since the scalability with the number of agents is often a computation bottleneck for game-theoretic planners, in Table 3, we further compare the computation speed of BRNE with two state-of-the-art dynamic game solvers, ALGAMES (Le Cleac’h et al., 2022) and iLQGames (Fridovich-Keil et al., 2020b). We use 100 samples per agent for BRNE and use the results reported in the “intersection” experiment from Le Cleac’h et al. (2022), as it is the most similar to the crowd navigation setting in this test. We use the importance-sampling-based BRNE implementation throughout this test. In Table 3, we can see that BRNE is at least one order of magnitude faster than ALGAMES and iLQGames and is the only method with sufficient computation time for real-time planning with 4 agents. ALGAMES and iLQGames are general-purpose game-theoretic planners that can be applied to applications other than crowd navigation, while BRNE is designed specifically for crowd navigation. BRNE explicitly benefits from its narrow scope for better computation efficiency in dense crowds.
@@
5.3 Simulated crowd navigation evaluation
@@
Table 5: Results of human dataset crowd navigation experiments.
@@
We select $3$ baselines to compare with: (1) ORCA; (2) CADRL; (3) Constrastive learning social navigation (NCE(Liu et al., 2021). We choose ORCA and CADRL for the same reason as we choose them for multi-agent navigation evaluation. We choose NCE since it is reported as the state-of-the-art method in simulated crowd navigation benchmarks. We train NCE in the environment provided by Liu et al. (2021).
@@ Results
In Table 4 and Figure 8, we show the safety and navigation efficiency results. In Figure 9, we show representative frames of our method from the simulated tests. We can see that our method BRNE is competitive on both navigation safety and efficiency with the state-of-the-art reinforcement learning method NCE. Both NCE and CADRL are tested in the same simulation environment where it is trained, while our method BRNE requires no training. CADRL performs poorly in this benchmark, exhibiting a high collision rate and low navigation efficiency. Furthermore, CADRL fails to reach the navigation goal and we have to terminate the trial after a certain amount of time, causing the low standard deviation in the time-to-goal metric. The performance of CADRL could be due to the difference between the testing environment provided by Liu et al. (2021) and the training environment of CADRL. Videos of the simulated crowd navigation experiments are included in our project website: https://sites.google.com/view/brne-crowdnav.
5.4 Crowd navigation in human datasets
@@
We evaluate crowd navigation with prerecorded human pedestrian behaviors within the same geometrical spaces of recordings. The evaluation is conducted in SocNavBench (Biswas et al., 2022), a state-of-the-art, high-fidelity crowd navigation benchmark framework. SocNavBench contains human pedestrian data from two of the most commonly used pedestrian datasets ETH (Pellegrini et al., 2009) and UCY (Lerner et al., 2007). A set of 33 curated episodes are extracted from the original datasets for benchmarking, which assemble a set of highly interactive test trials. The curated episodes contain an average of 44 pedestrians per episode. The distance between the goal location and the start location for the robot ensures that the navigation task can be finished within 25 seconds if the robot does not avoid collision and navigates with the maximum permitted velocity of $1.2m/s$. An occupancy grid map of the real-world space is provided for each episode, the robot has access to the map a priori and can only navigate in the free space. Note that, even though the pedestrian agents in SocNavBench are non-reactive given the pre-recorded human behavior, the robot agent still assumes the pedestrian agents are real-world reactive humans. We refer readers to Biswas et al. (2022) for more details about the benchmark, as well as the discussion on the choice of using pre-recorded human behavior. In this experiment, BRNE is implemented with importance sampling as a model predictive planner, updating its inference of mixed strategy Nash equilibrium based on the latest observation of other agents’ positions and velocities. The algorithm implementation for the benchmark (including paramerters) will be released on https://sites.google.com/view/brne-crowdnav.
@@
We compare our method BRNE with the four baselines and the corresponding results reported in the original paper of SocNavBench (Biswas et al., 2022). They are (1) the social force model (SF(Helbing and Molnár, 1995); (2) ORCA; (3) CADRL; and (4) a pedestrian-unaware baseline provided by the benchmark with static obstacle avoidance capability (Baseline). Note that ORCA and CADRL do not have static obstacle avoidance capability. Thus, they use Baseline as a meta-planner to avoid static obstacles in the benchmark. We tried to use Baseline directly as the meta-planner to generate the nominal mixed strategy for BRNE. However, even though Baseline is sufficient for BRNE to avoid static obstacles, it is not sufficient for BRNE to consistently reach the goal—recall that BRNE requires a strong goal-oriented nominal mixed strategy. Therefore, we slightly modified Baseline to keep its static obstacle avoidance module but improve its goal-reaching capability. We name this modified baseline planner Meta-Planner and report its performance as well. We will release the implementation of both BRNE and Meta-Planner in this benchmark.
@@ Results
Table 5 shows results of the human dataset crowd navigation experiments. Figure 10 shows representative snapshots of the benchmark tests. Both our method (BRNE) and the social force model (SF) significantly outperform the rest of the methods on navigation safety. Our method shares the minimal number of collisions with the social force model while being the only method with zero freezing behavior and exhibiting better path efficiency compared with the social force model. More importantly, compared to our meta-planner baseline Meta-planner(Ours), whose path is used by BRNE to generate the nominal mixed strategy, BRNE significantly improves navigation safety, reducing the number of pedestrian collisions from 37 to 1, while sharing nearly identical navigation efficiency as the baseline. As discussed in the next subsection, this one collision incident of our method is due to the pre-processing of the dataset instead of our method’s performance. These experimental results show that, when interacting with real human crowds in unstructured environments, our method can improve an existing navigation method’s safety to be near-perfect with minimal impact on navigation efficiency. This indicates that our method reaches human-level crowd navigation performance with a meta-planner on safety and navigation efficiency. The results also show great potential for deploying our method on other navigation pipelines. Videos of the human dataset experiments are included in our project website: https://sites.google.com/view/brne-crowdnav.
@@
While we will demonstrate our BRNE algorithm on robot hardware with real-world humans in the next section, we leave benchmarking our method in real-world crowds for future work, due to the time and financial cost of conducting such benchmarks. Here, we want to discuss the potential gap of sim-to-real transfer based on the human dataset benchmark results. First, since the robot is observing real-world human behavior in the benchmark, the benchmark results can at least verify which method will fail in the real world—a method that is dangerous or inefficient in the benchmark will likely remain dangerous or inefficient in the real world. Second, one main challenge of sim-to-real transfer in crowd navigation is the misalignment of human behavior assumptions. In our case, by inferring mixed strategy Nash equilibrium for navigation, the robot assumes the humans are cooperative, which is not necessarily true. However, the human dataset benchmark can serve as a preliminary verification of the influence of the misaligned human behavior assumption. Recall that, even though human agents are not reactive in the benchmark, the robot still believes it is interacting with real humans and observing real human behavior. The numerical efficiency of our method enables fast replanning during navigation, allowing the robot to constantly adjust its plan based on the latest observation, including observations that humans do not follow the assumed cooperative behavior. Furthermore, the robot makes navigation decisions based on mixed strategies—it does not anticipate exact human behavior but instead a distribution of likely human behaviors, which already considers unexpected human reactions. Thus, the fast replanning capability and mixed strategy-based modeling make the performance of our method in the human dataset benchmark more transferable in the real world. Lastly, our method is essentially built on top of an existing meta-planner provided by the benchmark. The modular design of our algorithm enhances the deployment flexibility, which lowers the gap of sim-to-real transfer. For example, our method can be integrated into existing robot navigation frameworks such as the navigation stack from the Robot Operating System (Macenski et al., 2020, 2022). This is particularly important for crowd navigation in unstructured real-world environments, and is not addressed by most existing crowd navigation and game-theoretic planning methods. Lastly, even though out of the scope of this paper, we want to point out one important aspect of the sim-to-transfer that is not addressed in the benchmark: the perception aspect of crowd navigation. The benchmark assumes near-perfect robot perception of pedestrian states, but robust perception for both robot localization and pedestrian tracking remains an open challenge and could influence the real-world deployment of our method.
5.5 Real-World Hardware Demonstration
@@
We implement our BRNE algorithm within the Isaac Robot Operation System (ROS) framework to communicate with the perception module and the robot’s low-level controller. We implement the algorithm in Python and used the PyTorch library to accelerate the weight updating step of the algorithm using the low-power GPU of the onboard computer. To maintain the real-time computation speed, we limited the maximum number of agents processed by the algorithm to 5, which allows the algorithm to control the robot at 10 Hz and replan at 5 Hz, with a planning horizon of 2 seconds and 200 trajectory samples per agent. We will release the hardware design and software implementation alongside the algorithm parameters. The algorithm implementation for the hardware demonstration (including parameters) will be released on https://sites.google.com/view/brne-crowdnav.
@@
The demonstrations are conducted in a $3m$ wide and $9m$ long space inside an indoor atrium. The demonstrations are designed to mimic the specifications of the crowd navigation tests with ORCA agents (such as the ones shown in Figure 9) and with pedestrian datasets (such as the ones shown in Figure 10). The number of pedestrians varies from $4$ to $10$. Our demonstrations follow common design principles in other crowd navigation works, such as Mavrogiannis et al. (2018) and Mavrogiannis et al. (2023a). To mimic the test with ORCA agents, pedestrians were instructed to go to the designated goals in a circle. To mimic the pedestrian dataset benchmark, pedestrians were instructed to move freely in the space. Across all demonstrations, the pedestrians are instructed to “move with normal walking speed and treat the robot as a walking person or dog”. The navigation task for the robot is to move safely and efficiently from one side of the room to the other, with a traveling distance varying between $6m$ to $8m$. The robot moves at a nominal speed of $0.5m/s$.
Discussion on Hardware Demonstration
The first challenge encountered during the hardware demonstration is specifying the risk function (Definition 6). We model the risk function as a logistic function based on the distance between two agents. The parameters of this function, which dictate how an agent’s willingness to be near others decreases with distance, are set manually through trial and error. Future work could involve learning the risk function from pedestrian datasets and dynamically adjusting its parameters based on real-time observations. A second challenge arises from inconsistent pedestrian observations due to occlusion. These inconsistencies cause agents to appear and disappear unexpectedly, leading to inaccuracies in pedestrian velocity estimation. To address this, we apply an iterative-closest-point method to associate pedestrians across frames, which reduces this issue but may still result in oscillations in velocity estimates, affecting the fidelity of nominal pedestrian strategies. Future work could explore data-driven filters to further stabilize pedestrian tracking.
6 Conclusion
@@
In this work, we propose a computation-efficient mixed strategy Nash equilibrium model for crowd navigation. Mixed strategy Nash equilibrium provides a high-level cooperation model for the robot to plan actions that leverage human cooperation for collision avoidance while maintaining the uncertainty in human behavior. Despite the general hardness of computing mixed strategy Nash equilibrium, we achieve real-time inference speed on a laptop CPU and a low-power embedded computer by establishing the formal connection between mixed strategy Nash equilibrium and a simple iterative Bayesian update scheme. We name the proposed model Bayesian Recursive Nash Equilibrium (BRNE). We further develop a model predictive crowd navigation framework using Gaussian processes to bootstrap agents’ nominal mixed strategies. The proposed model can be incorporated into existing navigation frameworks to navigate alongside humans in unstructured environments with static obstacles. Our experiment results show that our BRNE model significantly improves the safety of a human-unaware planner without compromising navigation efficiency in the human dataset benchmark, reaching human-level performance. Compared to other crowd navigation algorithms, our model consistently outperforms them in both safety and navigation efficiency. Lastly, we demonstrate the practicality of our model on an untethered robot quadruped for real-time crowd navigation with fully onboard perception and computation.